OpenStudy (anonymous):

Solve using the principle of zero products (13a+9) (4a-12) = 0

OpenStudy (saifoo.khan):

13a + 9 = 0 & 4a - 12 =0 13a = -9 & a = 12/4 a = -9/13 & a = 3

OpenStudy (anonymous):

hahaha i had to look up what principle of zero products was because i was taught it as a property not a principle